# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ 4.4 An Important Clarification: The Role of Purity
In complexity measures, the role played by purity, $P=\mathrm{Tr}(\rho^{2})$, is often misunderstood. A measurement of purity provides an upper bound on $S(\rho)$, not a lower bound, because the von Neumann entropy $S(\rho)$ is always greater than or equal to the second-order Rényi entropy $S_{2}(\rho)=-\log P$. Therefore, substituting this relationship into the expression for RCC, Eq. (2.3), a purity measurement can only provide an upper bound on complexity: $C_{R}(\rho)\leq\frac{(\log d_{R}+\log P)}{\log\Gamma_{R}}$. It cannot be used to prove a lower bound on complexity; its more appropriate role is as a sanity check tool. A correct understanding is that “high purity” in itself does not guarantee a large RCC lower bound. A meaningful lower bound must originate from the distinguishability of the state $\rho$ relative to the reference state $\sigma_{R}$. This distinguishability is quantified by the success rate of hypothesis testing, a high occupation probability on a low-rank projection, or a classical entropy that is far from the uniform distribution, rather than by purity itself. In contrast, the three operational principles proposed in this chapter can all provide rigorous lower bounds. Detailed protocols for the three principles mentioned above can be found in Appendix D, and their technical details and boundary conditions on issues such as multiple testing, statistical power control, and readout leakage propagation should be consulted in conjunction with Appendix C.
5 Properties and Application Interfaces of RCC
@@
The preceding sections have established the definition of Reference-Contingent Complexity (RCC), its theorem as a lower bound for optimal gate count, and the operational principles for extracting its information. A well-behaved physical quantity should not only be operational but also adhere to a series of universal physical principles and be able to seamlessly interface with existing theoretical frameworks. This chapter aims to reveal that RCC indeed possesses such qualities. We will first systematically elaborate on the axiomatic properties that RCC satisfies as a well-behaved resource measure, and subsequently showcase its three key application interfaces: first, we will elucidate how the introduction of RCC deepens and unifies the performance frontier of quantum dynamics; second, how it provides a structurally fair calibration scheme for exploring fundamental questions such as the complexity-geometry duality; and finally, this paper will reveal its third application as a tool for quantifying thermodynamic order and constraining statistical process laws, demonstrating how this state function can provide rigorous lower bounds for the information processing work and evolution time in physical processes.
5.1 Axiomatic Properties as a Resource Measure
@@
    Monotonicity under cluster-preserving channels: For a class of "free operations" that preserve the structure of the reference set [^2] Here, ”free operations” are formally defined as a class of cluster-preserving CPTP maps $E$ that maintain the structure of the reference set, with their core property being $E(\sigma_{R})=\sigma_{R}$. The monotonicity of RCC under this class of operations, $C_{R}(E(\rho))\leq C_{R}(\rho)$, is a direct corollary of the data processing inequality for quantum relative entropy, $D(E(\rho)\|E(\sigma_{R}))\leq D(\rho\|\sigma_{R})$.—that is, CPTP maps $E$ satisfying $E(\sigma_{R})=\sigma_{R}$—RCC is necessarily non-increasing:
    $C_{R}(E(\rho))\leq C_{R}(\rho)$ (5.1)
    This property originates from the data processing inequality for quantum relative entropy, $D(E(\rho)||E(\sigma_{R}))\leq D(\rho||\sigma_{R})$, and it ensures the self-consistency of RCC as a resource that cannot be created "for free," fully conforming to the general framework of quantum resource theories [13]. (Note: A rigorous definition of "free operations" and the objectivization criteria that should be followed for selecting a physically appropriate reference set are discussed in depth in Appendix F.3.)
@@
      a. If the hardware addressability remains unchanged ($\Gamma_{R^{\prime}}=\Gamma_{R}$), the complexity does not increase: $C_{R^{\prime}}(\rho)\leq C_{R}(\rho)$.
@@
      b. If the stronger constraint limits the addressability ($\Gamma_{R^{\prime}}<\Gamma_{R}$), then monotonicity is not guaranteed.
@@
5.2 Application Interface 1: Providing an Operational Lower Bound for the Dynamical Performance Equation
As an independent and self-consistent lower bound on complexity, an important application of RCC is to provide an operational interface for our previous work on the ’energy-entanglement performance equation’ [31][32, 33]. In that work, by unifying quantum speed limits and principles of locality, we established an "energy-entanglement performance equation" (RECT-$\eta$) that connects dynamical resources, information output, and the optimal computational complexity $C_{\mathrm{opt}}$. The core conclusion of that framework is a precise ’resource accounting identity’ (its full derivation is presented in Appendix E), which directly relates physical quantities of the system to theoretical complexity:
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (5.2)
where $\eta_{\mathrm{QSL}}$ and $\eta_{\mathrm{LR}}$ are dynamical factors that quantify the efficiency of the system’s utilization of energy resources and entanglement propagation, respectively; $\gamma$ is the locality constant of the LR bound (dimensionless); and $\sigma_{\mathrm{avail}}$ represents the available energy fluctuation / effective energy bandwidth (a process-dependent energy scale). Although Eq. (5.2) is theoretically universal, the $C_{\mathrm{opt}}$ it contains is a theoretical limit that cannot be directly measured, which restricts the direct experimental application of the framework. The RCC framework proposed in this work provides a solution to this. By applying our core result, Theorem 3.1, which provides a rigorous lower bound for $C_{\mathrm{opt}}$ with $C_{R}(\rho)$ (i.e., $C_{\mathrm{opt}}\geq C_{R}(\rho)$), we can transform the above theoretical identity into a performance inequality that can be directly used for experimental and numerical analysis:
$\sigma_{\mathrm{avail}}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{R}(\rho)$ (5.3)
@@ 5.2 Application Interface 1: Providing an Operational Lower Bound for the Dynamical Performance Equation
To obtain a more universal, dimensionless expression that facilitates cross-platform comparisons, we divide both sides of Eq. (5.3) by the system’s characteristic energy scale $J$, yielding:
$\frac{\sigma_{\mathrm{avail}}}{J}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma}{2}\right)\cdot C_{R}(\rho)$ (5.4)
Eq. (5.4) represents a significant conceptual deepening. The introduction of RCC transforms the core complexity of the original framework from an abstract quantity $C_{\mathrm{opt}}$, which is dependent on the specific evolution path and is not measurable, into a state function $C_{R}$, which is determined solely by the final state $\rho$ of the system and is, in principle, measurable. This reveals a more profound and more operational physical constraint: regardless of the ingenious path the system evolves through, the final achievable "resource-output product" (the product of energy fluctuation normalized by the characteristic energy scale and the final state’s entanglement entropy) is strictly lower-bounded by the structurally fair complexity $C_{R}$ inherent in the final state itself. Therefore, RCC plays a key bridging role here, transforming a purely theoretical relation (the RECT-$\eta$ identity) into a falsifiable physical proposition directly related to the properties of the final state, thereby greatly expanding the applicability of the original framework. The complete derivation of the aforementioned performance equation, including its profound physical meaning as a "resource accounting identity," is detailed in Appendix E.
5.3 Application Interface 2: Providing a Calibration Scheme for the Complexity-Geometry Duality
In frontier fields such as holographic gravity, conjectures like "complexity = geometry" (e.g., complexity = volume or complexity = action) are highly inspiring, but one of their core obstacles is that the definition of complexity on the boundary field theory side is ambiguous and non-unique [34]. The "structural fairness" idea of RCC provides a "bottom-up," physically well-defined proposal for a structured benchmark. We propose a testable calibration procedure to explore this dual relationship:
@@ 5.3 Application Interface 2: Providing a Calibration Scheme for the Complexity-Geometry Duality
5.4 Application Interface 3: Quantifying Thermodynamic Order and Constraining Process Laws
Following the preceding interfaces that connect RCC with the system’s dynamics (Sec. 5.2) and spacetime geometry (Sec. 5.3), this section aims to reveal its third application: as a tool for quantifying thermodynamic order and constraining statistical process laws. A central challenge in physics is how to use easily measurable "state functions" (such as entropy, free energy) to constrain hard-to-track "process quantities" (such as work, heat, evolution time). This interface aims to provide a novel solution to this challenge. We will demonstrate that two core thermo-dynamical resources consumed by a quantum process, Information Processing Work and Evolution Time, can be rigorously lower-bounded by an operational state function determined solely by the initial and final states of the system, namely, Reference-Contingent Complexity (RCC). All technical derivation details are consolidated in Appendix I. To achieve the above goal, we leverage the theoretical framework of "Complexity Window Thermodynamics" (CWT) [35]. CWT generalizes statistical mechanics to handle information complexity by introducing a finite "complexity budget" $\Xi$. Its core physical quantities include the windowed entropy $S_{\Xi}$, effective temperature $T_{\Xi}$, and information processing work $\Pi_{\Xi}$. To work within this framework, we need to naturally generalize RCC to a quantity that depends on the observational capability $\Xi$. In the following, the partial derivative with respect to $\Xi$, unless otherwise specified, is assumed to be taken at a fixed energy constraint $E$ and is denoted as $(\partial/\partial\Xi)_{E}$.
@@ 5.4 Application Interface 3: Quantifying Thermodynamic Order and Constraining Process Laws
$C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (5.5)
where $S_{\Xi}(\rho)=S(E_{\Xi}(\rho))$ is the windowed entropy of $\rho$ under the conditional expectation map $E_{\Xi}$, and $S$ is the von Neumann entropy. Physically, $C_{R,\Xi}(\rho)$ quantifies the non-trivial degree of order exhibited by the system $\rho$ beyond its a priori structure, under a finite observational capability $\Xi$. As proven in Appendix I.2, $S_{\Xi}$ is monotonically non-increasing with $\Xi$, and therefore $C_{R,\Xi}$ is monotonically non-decreasing with $\Xi$; in the limit $\Xi\to\infty$, it recovers the standard RCC of Sec. 2.3. This generalization is central to connecting the two theoretical frameworks, as it allows us to relate the process quantities in CWT to the state function RCC.
@@
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (5.6)
Proof Outline. The result is obtained by directly differentiating the definition of $C_{R,\Xi}$ and combining it with the definition of $\Pi_{\Xi}$ in CWT; its rigorous derivation, including the more general case of a variable bandwidth $\Gamma_{R}=\Gamma_{R}(\Xi)$, is detailed in Appendix I.3. This identity serves as a bridge between thermodynamics and information theory: the energy cost to increase the system’s resolvable complexity ($\Pi_{\Xi}$) is proportional to the product of the system’s energy scale ($T_{\Xi}$) and the complexity gain with an increasing budget $($$\partial C_{R,\Xi}/\partial\Xi)_{E}$), with the proportionality constant being the hardware’s information bandwidth ($\log\Gamma_{R}$). Rigorously, the total information processing work is $W_{\mathrm{info}}=\int\Pi_{\Xi}\mathrm{d}\Xi=\log\Gamma_{R}\cdot\int T_{\Xi}\mathrm{d}C_{R,\Xi}$. When $T_{\Xi}$ varies slowly within the interval, $W_{\mathrm{info}}\approx\bar{T}_{\Xi}\cdot\log\Gamma_{R}\cdot\Delta C_{R,\Xi}$, where $\bar{T}_{\Xi}$ is the weighted average temperature along the complexity coordinate (its rigorous definition is given in Appendix I.3.1). This result directly links the macroscopic process quantity $W_{\mathrm{info}}$ with the net change in a state function, $\Delta C_{R,\Xi}$, which is determined solely by the information structure of the initial and final states, providing a generalization of Landauer’s principle [36] for "structured information." In addition to constraining the energy cost, RCC also sets a fundamental lower limit on the time cost of a quantum process.
Theorem 5.4(RCC-Driven Lower Bound on Process Time). Consider a quantum process evolving from an initial state $\rho_{i}$ to a final state $\rho_{f}$ over a total duration $t_{f}$, with an initial optimal complexity $C^{*}(0)$. If the optimal complexity evolution envelope $C^{*}(t)$ of the process adheres to the universal rate upper bound of CWT, $\dot{C}^{*}\leq(2/\hbar)\Pi_{\Xi}$ [35] Prop. 6.1, then
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (5.7)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (5.8)
@@
If the process is approximately isothermal, and one adopts the chaos-bound analogy from CWT, $\Pi_{\Xi}\lesssim\pi k_{B}T_{\Xi}$, then $t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$, which formally resonates with the chaos time scale bound of Maldacena et al. [37]. This result is complementary to the RECT-$\eta$ framework of Sec. 5.2, indicating that the time cost of a quantum process is ultimately constrained by the information-theoretic complexity it must generate. The introduction of RCC allows us to establish a bridge, unified by "information structure complexity," across the three major physical domains of dynamics (Sec. 5.2), geometry (Sec. 5.3), and thermodynamic statistics (this section). The RCC terms in Eqs. (5.7) and (5.8) can, in principle, be experimentally estimated via the operational protocols proposed in Chapter 4; a robustness framework for handling real-world non-idealities (such as state leakage and statistical errors) is provided in Appendix I.5. This paves the way for experimentally testing these fundamental "time-work-information" trade-off relations in real physical systems, and suggests that the computational complexity of a quantum state may not be merely a metric in computational science, but rather a deeper organizing principle of the material world that unifies its various physical manifestations.
6 Discussion

### v3

@@ IV.4 An Important Clarification: The Role of Purity
In complexity measures, the role played by purity, $P=\mathrm{Tr}(\rho^{2})$, is often misunderstood. A measurement of purity provides an upper bound on $S(\rho)$, not a lower bound, because the von Neumann entropy $S(\rho)$ is always greater than or equal to the second-order Rényi entropy $S_{2}(\rho)=-\log P$. Therefore, substituting this relationship into the expression for RCC, Eq. (3), a purity measurement can only provide an upper bound on complexity: $C_{R}(\rho)\leq\frac{(\log d_{R}+\log P)}{\log\Gamma_{R}}$. It cannot be used to prove a lower bound on complexity; its more appropriate role is as a sanity check tool. A correct understanding is that “high purity” in itself does not guarantee a large RCC lower bound. A meaningful lower bound must originate from the distinguishability of the state $\rho$ relative to the reference state $\sigma_{R}$. This distinguishability is quantified by the success rate of hypothesis testing, a high occupation probability on a low-rank projection, or a classical entropy that is far from the uniform distribution, rather than by purity itself. In contrast, the three operational principles proposed in this chapter can all provide rigorous lower bounds. Detailed protocols for the three principles mentioned above can be found in Appendix D, and their technical details and boundary conditions on issues such as multiple testing, statistical power control, and readout leakage propagation should be consulted in conjunction with Appendix C.
V Properties and Application Interfaces of RCC
@@
The preceding sections have established the definition of Reference-Contingent Complexity (RCC), its theorem as a lower bound for optimal gate count, and the operational principles for extracting its information. A consistent resource measure is expected to satisfy universal physical axioms (such as monotonicity) and be compatible with established theoretical frameworks. This chapter aims to reveal that RCC indeed possesses such qualities. We will first systematically elaborate on the axiomatic properties that RCC satisfies as a well-behaved resource measure, and subsequently showcase its three key application interfaces: first, we will elucidate how the introduction of RCC deepens and unifies the performance frontier of quantum dynamics; second, how it provides a structurally fair calibration scheme for exploring fundamental questions such as the complexity-geometry duality; and finally, this paper will reveal its third application as a tool for quantifying thermodynamic order and constraining statistical process laws, demonstrating how this state function can provide rigorous lower bounds for the information processing work and evolution time in physical processes.
V.1 Axiomatic Properties as a Resource Measure
@@
    Monotonicity under cluster-preserving channels: For a class of "free operations" that preserve the structure of the reference set [^2] Here, “free operations” are formally defined as a class of cluster-preserving CPTP maps $E$ that maintain the structure of the reference set, with their core property being $E(\sigma_{R})=\sigma_{R}$. The monotonicity of RCC under this class of operations, $C_{R}(E(\rho))\leq C_{R}(\rho)$, is a direct corollary of the data processing inequality for quantum relative entropy, $D(E(\rho)\|E(\sigma_{R}))\leq D(\rho\|\sigma_{R})$.—that is, CPTP maps $E$ satisfying $E(\sigma_{R})=\sigma_{R}$—RCC is necessarily non-increasing:
    $C_{R}(E(\rho))\leq C_{R}(\rho)$ (12)
    This property originates from the data processing inequality for quantum relative entropy, $D(E(\rho)||E(\sigma_{R}))\leq D(\rho||\sigma_{R})$, and it ensures the self-consistency of RCC as a resource that cannot be created "for free," fully conforming to the general framework of quantum resource theories [BrandaoGour2015]. (Note: A rigorous definition of "free operations" and the objectivization criteria that should be followed for selecting a physically appropriate reference set are discussed in depth in Appendix F.3.)
@@
      If the hardware addressability remains unchanged ($\Gamma_{R^{\prime}}=\Gamma_{R}$), the complexity does not increase: $C_{R^{\prime}}(\rho)\leq C_{R}(\rho)$.
@@
      If the stronger constraint limits the addressability ($\Gamma_{R^{\prime}}<\Gamma_{R}$), then monotonicity is not guaranteed.
@@
V.2 Application Interface 1: Providing an Operational Lower Bound for the Dynamical Performance Equation
As an independent and self-consistent lower bound on complexity, an important application of RCC is to provide an operational interface for our previous work on the ’energy-entanglement performance equation’ [liu2025cost][MandelstamTamm1945, LiebRobinson1972]. In that work, by unifying quantum speed limits and principles of locality, we established an "energy-entanglement performance equation" (RECT-$\eta$) that connects dynamical resources, information output, and the optimal computational complexity $C_{\mathrm{opt}}$. The core conclusion of that framework is a precise ’resource accounting identity’ (its full derivation is presented in Appendix E), which directly relates physical quantities of the system to theoretical complexity:
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (13)
where $\eta_{\mathrm{QSL}}$ and $\eta_{\mathrm{LR}}$ are dynamical factors that quantify the efficiency of the system’s utilization of energy resources and entanglement propagation, respectively; $\gamma$ is the locality constant of the LR bound (dimensionless); and $\sigma_{\mathrm{avail}}$ represents the available energy fluctuation / effective energy bandwidth (a process-dependent energy scale). Although Eq. (13) is theoretically universal, the $C_{\mathrm{opt}}$ it contains is a theoretical limit that cannot be directly measured, which restricts the direct experimental application of the framework. The RCC framework proposed in this work provides a solution to this. By applying our core result, Theorem 3.1, which provides a rigorous lower bound for $C_{\mathrm{opt}}$ with $C_{R}(\rho)$ (i.e., $C_{\mathrm{opt}}\geq C_{R}(\rho)$), we can transform the above theoretical identity into a performance inequality that can be directly used for experimental and numerical analysis:
$\sigma_{\mathrm{avail}}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{R}(\rho)$ (14)
@@ V.2 Application Interface 1: Providing an Operational Lower Bound for the Dynamical Performance Equation
To obtain a more universal, dimensionless expression that facilitates cross-platform comparisons, we divide both sides of Eq. (14) by the system’s characteristic energy scale $J$, yielding:
$\frac{\sigma_{\mathrm{avail}}}{J}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma}{2}\right)\cdot C_{R}(\rho)$ (15)
Eq. (15) reformulates the dynamical constraint in terms of state complexity. The introduction of RCC transforms the core complexity of the original framework from an abstract quantity $C_{\mathrm{opt}}$, which is dependent on the specific evolution path and is not measurable, into a state function $C_{R}$, which is determined solely by the final state $\rho$ of the system and is, in principle, measurable. This reveals a more profound and more operational physical constraint: regardless of the ingenious path the system evolves through, the final achievable "resource-output product" (the product of energy fluctuation normalized by the characteristic energy scale and the final state’s entanglement entropy) is strictly lower-bounded by the structurally fair complexity $C_{R}$ inherent in the final state itself. Therefore, RCC plays a key bridging role here, transforming a purely theoretical relation (the RECT-$\eta$ identity) into a falsifiable physical proposition directly related to the properties of the final state, thereby greatly expanding the applicability of the original framework. The complete derivation of the aforementioned performance equation, including its profound physical meaning as a "resource accounting identity," is detailed in Appendix E.
V.3 Application Interface 2: Providing a Calibration Scheme for the Complexity-Geometry Duality
In frontier fields such as holographic gravity, conjectures like "complexity = geometry" (e.g., complexity = volume or complexity = action) are highly inspiring, but one of their core obstacles is that the definition of complexity on the boundary field theory side is ambiguous and non-unique [Brown2016CEA]. The "structural fairness" idea of RCC provides a "bottom-up," physically well-defined proposal for a structured benchmark. We propose a testable calibration procedure to explore this dual relationship:
@@ V.3 Application Interface 2: Providing a Calibration Scheme for the Complexity-Geometry Duality
V.4 Application Interface 3: Quantifying Thermodynamic Order and Constraining Process Laws
Following the preceding interfaces that connect RCC with the system’s dynamics (Sec. V.2) and spacetime geometry (Sec. V.3), this section aims to reveal its third application: as a tool for quantifying thermodynamic order and constraining statistical process laws. A central challenge in physics is how to use easily measurable "state functions" (such as entropy, free energy) to constrain hard-to-track "process quantities" (such as work, heat, evolution time). This interface aims to provide a novel solution to this challenge. We will demonstrate that two core thermo-dynamical resources consumed by a quantum process, Information Processing Work and Evolution Time, can be rigorously lower-bounded by an operational state function determined solely by the initial and final states of the system, namely, Reference-Contingent Complexity (RCC). All technical derivation details are consolidated in Appendix I. To achieve the above goal, we leverage the theoretical framework of "Complexity Window Thermodynamics" (CWT) [liu2025heuristic]. CWT generalizes statistical mechanics to handle information complexity by introducing a finite "complexity budget" $\Xi$. Its core physical quantities include the windowed entropy $S_{\Xi}$, effective temperature $T_{\Xi}$, and information processing work $\Pi_{\Xi}$. To work within this framework, we need to naturally generalize RCC to a quantity that depends on the observational capability $\Xi$. In the following, the partial derivative with respect to $\Xi$, unless otherwise specified, is assumed to be taken at a fixed energy constraint $E$ and is denoted as $(\partial/\partial\Xi)_{E}$.
@@ V.4 Application Interface 3: Quantifying Thermodynamic Order and Constraining Process Laws
$C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (16)
where $S_{\Xi}(\rho)=S(E_{\Xi}(\rho))$ is the windowed entropy of $\rho$ under the conditional expectation map $E_{\Xi}$, and $S$ is the von Neumann entropy. Physically, $C_{R,\Xi}(\rho)$ quantifies the non-trivial degree of order exhibited by the system $\rho$ beyond its a priori structure, under a finite observational capability $\Xi$. As proven in Appendix I.2, $S_{\Xi}$ is monotonically non-increasing with $\Xi$, and therefore $C_{R,\Xi}$ is monotonically non-decreasing with $\Xi$; in the limit $\Xi\to\infty$, it recovers the standard RCC of Sec. II.3. This generalization is central to connecting the two theoretical frameworks, as it allows us to relate the process quantities in CWT to the state function RCC.
@@
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (17)
Proof Outline. The result is obtained by directly differentiating the definition of $C_{R,\Xi}$ and combining it with the definition of $\Pi_{\Xi}$ in CWT; its rigorous derivation, including the more general case of a variable bandwidth $\Gamma_{R}=\Gamma_{R}(\Xi)$, is detailed in Appendix I.3. This identity serves as a bridge between thermodynamics and information theory: the energy cost to increase the system’s resolvable complexity ($\Pi_{\Xi}$) is proportional to the product of the system’s energy scale ($T_{\Xi}$) and the complexity gain with an increasing budget $($$\partial C_{R,\Xi}/\partial\Xi)_{E}$), with the proportionality constant being the hardware’s information bandwidth ($\log\Gamma_{R}$). Rigorously, the total information processing work is $W_{\mathrm{info}}=\int\Pi_{\Xi}\mathrm{d}\Xi=\log\Gamma_{R}\cdot\int T_{\Xi}\mathrm{d}C_{R,\Xi}$. When $T_{\Xi}$ varies slowly within the interval, $W_{\mathrm{info}}\approx\bar{T}_{\Xi}\cdot\log\Gamma_{R}\cdot\Delta C_{R,\Xi}$, where $\bar{T}_{\Xi}$ is the weighted average temperature along the complexity coordinate (its rigorous definition is given in Appendix I.3.1). This result directly links the macroscopic process quantity $W_{\mathrm{info}}$ with the net change in a state function, $\Delta C_{R,\Xi}$, which is determined solely by the information structure of the initial and final states, providing a generalization of Landauer’s principle [Landauer_Ref] for "structured information." In addition to constraining the energy cost, RCC also sets a fundamental lower limit on the time cost of a quantum process.
Theorem 5.4(RCC-Driven Lower Bound on Process Time). Consider a quantum process evolving from an initial state $\rho_{i}$ to a final state $\rho_{f}$ over a total duration $t_{f}$, with an initial optimal complexity $C^{*}(0)$. If the optimal complexity evolution envelope $C^{*}(t)$ of the process adheres to the universal rate upper bound of CWT, $\dot{C}^{*}\leq(2/\hbar)\Pi_{\Xi}$ [liu2025heuristic] Prop. 6.1, then
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (18)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (19)
@@
If the process is approximately isothermal, and one adopts the chaos-bound analogy from CWT, $\Pi_{\Xi}\lesssim\pi k_{B}T_{\Xi}$, then $t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$, which formally resonates with the chaos time scale bound of Maldacena et al. [MSS2016]. This result is complementary to the RECT-$\eta$ framework of Sec. V.2, indicating that the time cost of a quantum process is ultimately constrained by the information-theoretic complexity it must generate. The introduction of RCC allows us to establish a bridge, unified by "information structure complexity," across the three major physical domains of dynamics (Sec. V.2), geometry (Sec. V.3), and thermodynamic statistics (this section). The RCC terms in Eqs. (18) and (19) can, in principle, be experimentally estimated via the operational protocols proposed Sec. IV; a robustness framework for handling real-world non-idealities (such as state leakage and statistical errors) is provided in Appendix I.5. This paves the way for experimentally testing these fundamental "time-work-information" trade-off relations in real physical systems, and suggests that the computational complexity of a quantum state may not be merely a metric in computational science, but rather a deeper organizing principle of the material world that unifies its various physical manifestations.
VI Discussion